find the value of 1/2sec80degree-2cos20degree.

Grade:11

1 Answers

Atish Dutta
8 years ago
1/2sec80degree-2cos20degree
=1/2cos80degree-2cos20degree
=1-2.2cos80degreecos20degree/2cos80degree
=1-2(cos100degree+cos60degree)/2cos80degree
=[1-2cos(180degree-80degree)-2.1/2]/2cos80degree
=1